a random sample of 81 items is taken, producing a sample mean of 47. The population standard deviation is 5.89. construct a 90 confidence interval to estimate the population mean.



1
Expert's answer
2022-06-09T12:11:04-0400

The critical value for "\alpha = 0.1" is "z_c = z_{1-\alpha/2} = 1.6449."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-z_c\times\dfrac{\sigma}{\sqrt{n}},\bar{x}+z_c\times\dfrac{\sigma}{\sqrt{n}})"

"=(47-1.6449\times\dfrac{5.89}{\sqrt{81}},47+1.6449\times\dfrac{5.89}{\sqrt{81}})"

"=(45.9235, 48.0765)"

Therefore, based on the data provided, the 90% confidence interval for the population mean is "45.9235 < \mu < 48.0765

," which indicates that we are 90% confident that the true population mean "\mu" is contained by the interval "(45.9235, 48.0765)."